OpenStudy (anonymous):

if the roots of the equation (a2+b2)x2-2(ac+bd)x+(c2+d2)=0 are equal, prove that a/b=c/d

OpenStudy (anonymous):

if two roots are equal, b2-4ac=0, or (-2(ac+bd))^2 - 4 (a^2 + b^2) (c^2+d^2) = 0 or 4(ac)^2 + 8abcd + 4(bd)^2 - 4(ac)^2 - 4(ad)^2 - 4(bc)^2 - 4(bd)^2 = 0 or 8abcd -4(ad)^2 -4(bc)^2 = 0 or (ad)^2 - 2(ad)(bc) + (bc)^2 = 0 or (ad - bc)^2 = 0 or ad - bc = 0 or ad = bc therefore, a/b = c/d

OpenStudy (anonymous):

thank you very much u r ginus

OpenStudy (anonymous):

u dont need to say that...........just click on GOOD ANSWER that is ok...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

only one more question please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

find the are of circle x2+y2=a2 by integration

OpenStudy (anonymous):

the answer is pi a^2 i have no idea if we can solve with integration........

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u mail me on devil.magra@yahoo.com becoz i will loss my train

OpenStudy (anonymous):

ok